OpenStudy (niahloves):

Which equation with exponential expressions is true? A. 33 = 3 x 3 B. 54 = 4 x 4 x 4 x 4 C. 67 = 7 x 7 x 7 x 7 x 7 x 7 D. 52 = 5 x 5

OpenStudy (niahloves):

@AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

what you think?

OpenStudy (niahloves):

i don't know??

OpenStudy (alexandervonhumboldt2):

none of them........

OpenStudy (niahloves):

what?

OpenStudy (alexandervonhumboldt2):

@ganeshie8 misterious question

OpenStudy (anonymous):

none of them seem to be right

OpenStudy (niahloves):

but you have to pick one

OpenStudy (anonymous):

did u right d correctly

OpenStudy (alexandervonhumboldt2):

not seems to be @YK.PAPI they are incorrect

OpenStudy (anonymous):

did she type d correctly

OpenStudy (niahloves):

no sorry here yall go

OpenStudy (alexandervonhumboldt2):

ohhhhhhhhh

OpenStudy (alexandervonhumboldt2):

a^2=a*a hint

OpenStudy (niahloves):

I think it's c or d

OpenStudy (alexandervonhumboldt2):

D because 5^2=5*5

OpenStudy (anonymous):

i say d

OpenStudy (alexandervonhumboldt2):

using the rule a^2=a*a a^3=a*a*a a^4=a*a*a*a and so on

OpenStudy (niahloves):

yeah it's d becuase C is wrong becuase you suppose to do 6 x6 seven times

OpenStudy (alexandervonhumboldt2):

yeah

OpenStudy (anonymous):

exactly